ROTATIONAL SPECTRA AND STRUCTURE OF WEAKLY BOUND Ar(H_{2}S)_{2} AND Ar(D_{2}S)_{2} COM-PLEXES

P. K. MANDAL, \underline{MAUSUMI GOSWAMI}, E. ARUNAN, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore 560 012 INDIA.

Rotational spectra of Ar(H_{2}S)_{2} and Ar(D_{2}S)_{2} complexes have been observed with a pulsed nozzle Fourier transform microwave spectrometer. About 40 a and b dipole transitions have been observed for each isotopomer. Each transition is spilt in to two as observed earlier by Lovas for the (H_{2}S)_{2} complex ^{a}. The rotational constants for the lower state of Ar(H_{2}S)_{2} are: A = 1810.410(6) MHz; B = 1596.199(9) MHz and C = 848.814(2) MHz; and those for Ar(D_{2}S)_{2} are: A = 1725.49(1) MHz, B = 1566.27(3) MHz and C = 826.817(4) MHz. The C rotational constants for the two states were nearly identical, for both isotopomers. The A and B rotational constants for the upper state of Ar(H_{2}S)_{2} are about 10 MHz and 6 MHz larger than those for the lower state. However, for Ar(D_{2}S)_{2}, A and B rotational constants for the upper state were larger only by 30 kHz and 50 kHz, respectively. This is in contrast to the very similar splittings observed in B rotational constants for (H_{2}S)_{2} (1.2 MHz) and (D_{2}S)_{2} (0.9 MHz). Assuming H_{2}S to be a sphere, the c.m. separation between two H_{2}S units is calculated to be 4.03 Åwhich is 0.1 Åless than that found in (H_{2}S)_{2} dimer. The distance between Ar and c.m. of (H_{2}S)_{2} is 3.55 Åand the Ar-c.m.(H_{2}S) distance is 4.09 Å. Ab initio calculations at MP2 level using various basis sets lead to three distinct minima including a pseudo-linear local minimum. At MP2/6-311++G(3df,2p) level of theory, the global minimum has a structure having Ar along the ’c’ axis of (H_{2}S)_{2}.

aF. J. Lovas private communication